# Workshop on Mechanistic Interpretability

### v4

@@ 1 Introduction
Retrieval-Augmented Generation (RAG), leveraging large language models (LLMs), has demonstrated significant potential in both academic research (qian2024grounding, ; yue2025inference, ; song2025measuring, ) and industrial applications (yang2024qwen2, ; guo2025deepseek, ) by enhancing the accuracy and grounding of generated responses through external contexts such as provided documents or retrieved articles online. A key benefit of RAG lies in its ability to mitigate the hallucination by explicitly attributing generated responses to specific segments of the provided context, known as context attribution [^1] We use the term context attribution in this work, and there are several different terms used in this area, such as citation, self-citation, etc. (wang2024backtracing, ; qi2024model, ; cohen2024contextcite, ; chuang2025selfcite, ).
@@
Nevertheless, verifying the extent to which generated responses are genuinely grounded in their cited context remains a challenging task. Current approaches frequently rely heavily on human annotation (menick2022teaching, ; slobodkin2024attribute, ) or computationally expensive methods such as model fine-tuning and gradient-based feature attribution for accurate attribution (yue2023automatic, ; qi2024model, ; chuang2025selfcite, ), particularly when dealing with extensive documents. For instance, qi2024model utilised distribution shifts between responses generated with and without context to identify relevant tokens and employed gradient-based feature attribution to pinpoint context relevance. Similarly, chuang2025selfcite enhanced context attribution accuracy through reward-driven fine-tuning within a Direct Preference Optimisation (DPO) framework, based on probability drop and hold analysis of model outputs to context ablation.
@@
To circumvent these computationally intensive methods, cohen2024contextcite introduced an inference-time attribution mechanism premised on the assumption that if removing grounded context segments substantially reduces the probability of a generated response, those segments are deemed necessary. Conversely, if retaining only grounded segments maintains response probability, these segments are considered sufficient. By capturing hundreds of probability ablation variations per context-response pair, cohen2024contextcite trained a linear surrogate model based on those hundreds of vectors, including the context segment masks and the corresponding generation probability of the original response, to identify context segments crucial for grounding model responses.
@@
However, cohen2024contextcite still need hundreds of RAG model’s forward calls to collect probability ablation samples for the linear surrogate model training. We propose a novel inference-time Jensen–Shannon Divergence driven method to Attribute Response to Context (ARC-JSD), building upon the inference-attribution assumption above. Our method evaluates the divergence in response distributions generated under the full context compared to sentence-ablated contexts, ranking context sentences based on their JSD differences because of JSD’s symmetric, finite, scale-free, and bounded properties (see § 3 and § 8 for details). This approach offers a significant computational advantage, as it eliminates the need for any additional fine-tuning or surrogate modelling. Furthermore, our ARC-JSD can avoid missing or smoothing non-linearities using JSD to directly quantify actual output distribution shift compared to the linear surrogate modelling (cohen2024contextcite, ).
@@
We empirically evaluate our JSD-driven context attribution approach across multiple question-answering benchmarks, i.e., TyDi QA (clark2020tydi, ), Hotpot QA (yang2018hotpotqa, ), and MuSiQue (trivedi2022musique, ), using state-of-the-art instruction-tuned LLMs including Qwen2-1.5B-Instruct, Qwen2-7B-Instruct (yang2024qwen2, ), Gemma2-2B-Instruct, and Gemma2-9B-Instruct (team2024gemma, ). Our results not only demonstrate improved average accuracy over 10% in context attribution but also achieve computational efficiency, achieving up to a three-fold speedup compared to cohen2024contextcite ’s linear-surrogate-based and other gradient-based baselines.
@@
Moreover, we investigate deeper into a mechanistic exploration of context attribution within RAGs by integrating JSD-based analysis with Logit Lens (Nostalgebraist2020, ). Through systematic probing, we identify specific attention heads and MLP layers critical for context attribution. By subsequently manipulating these located attention heads and MLPs using JSD scores as confidence gates, we can control RAGs and further mitigate the hallucination rate. In addition, we can visualise how relevant knowledge is stored in the corresponding MLP layers.
@@ 2 Related Work
Prior works for context attribution mainly focus on teaching RAG LLMs to generate self-citations for responses, such as few-shot in-context learning (gao2023enabling, ), instruction fine-tuning (ye2024effective, ). Some post-hoc works (chen2023purr, ; qi2024model, ) used an auxiliary language model or gradient-based feature attribution to locate relevant context segments. In general, those methods for context attribution are corroborative (worledge2024unifying, ) in nature, as citations within context are evaluated on whether they support or imply a generated response. Meanwhile, cohen2024contextcite ; chuang2025selfcite ; cohen2025learning ; liu2024attribot including our work focus on the contributive attribution methods, which are used to identify whether citations cause RAG LLMs to generate a response. chuang2025selfcite proposed a reward-based fine-tuning with DPO to guide RAG LLMs for context attribution, and cohen2024contextcite ; cohen2025learning further trained a linear surrogate model to identify context segments crucial for grounding model responses. liu2024attribot focuses on formalising and comparing different attribution acceleration methods and ignores attribution accuracy improvements. However, compared to cohen2024contextcite ; chuang2025selfcite ; cohen2025learning and corroborative methods above, our ARC-JSD method eliminates the need for any additional fine-tuning or surrogate modelling, and it can be directly integrated into any existing RAGs.
Mechanistic analysis for RAG.
@@
Existing mechanistic studies focus on next token generation task to analyse internal mechanisms of attention heads or MLPs, such as hallucination detection (ferrando2025do, ), multiple-choice questions (li2024anchored, ; wiegreffe2025answer, ; wang2025eliminating, ) and knowledge editing (meng2022locating, ; meng2023massediting, ; katz2024backward, ). Recently, sun2025redeep used a mechanistic interpretability method to analyse attention heads and MLPs of RAGs for the hallucination detection task. Compared to sun2025redeep focusing on locating sources which leads to hallucinations, our proposed ARC-JSD can be regarded as a complementary method to locate citations within context segments and analyse attentions and MLPs, which causes RAG LLMs to generate a correct response. wu2025retrieval focuses on mechanistically analysing retrieval attention heads of RAG LLMs under the Needle-in-the-Haystack (NIAH) setting, where they mainly evaluate whether retrieval attention heads conduct a copy-and-paste operation for retrieving a semantically irrelevant “needle” sentence from the context to the model’s outputs. Compared to wu2025retrieval , which restricts their mechanistic analysis to the NIAH setting where the model performs copy-and-paste retrieval, our work investigates how RAG LLMs mechanistically generate responses based on retrieved content through paraphrasing and contextual integration. This setting better reflects real-world RAG applications [^2] Compared to traditional RAG to directly map prompts based on their word embeddings, our work has a more general setting, which avoids potential embedding mismatch due to the common paraphrase of RAGs., where models rarely copy text exactly but instead synthesise and rephrase information from retrieved sources.
@@ 3 Background
Consider an autoregressive Transformer-based language model (LLM), denoted as $\mathcal{P}_{\text{LM}}(\cdot)$. Under RAG settings, this model generates responses ($\mathcal{R}$) based on an input query ($\mathcal{Q}$) and associated context ($\mathcal{C}$). Formally, response generation process can be described as $\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|\mathcal{C},\mathcal{Q})$, where context $\mathcal{C}$ consists of sentences $(c_{1},c_{2},\dots,c_{|\mathcal{C}|})$, the query $\mathcal{Q}$ comprises tokens $(q_{1},q_{2},\dots,q_{|\mathcal{Q}|})$, and the generated response $\mathcal{R}$ includes tokens $(r_{1},r_{2},\dots,r_{|\mathcal{R}|})$. Our analysis of context attribution focuses on how the entire response distribution changes when conditioned on the full context set and ablated context alongside the query: $\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|c_{1},\dots,c_{|\mathcal{C}|},\mathcal{Q}),\>\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|\mathcal{C}_{\text{ABLATE}}(c_{i}),\mathcal{Q})\,\text{where}\,\mathcal{C}_{\text{ABLATE}}(c_{i})=\mathcal{C}\setminus\{c_{i}\},\>i\in\{1,\dots,|\mathcal{C}|\}$.
Logit Lens.
@@
Logit lens (Nostalgebraist2020, ) is a mechanistic interpretability method designed to analyse intermediate representations within autoregressive Transformers. Given the LLM architecture described in Appendix D, logit lens leverages intermediate representations to quantify the direct contribution of attention heads ($\mathbf{a}_{i}^{\ell,h}$), MLP outputs ($\mathbf{m}_{i}^{\ell}$), and residual streams ($\mathbf{x}_{i}^{\ell}$) to token logits: $\text{logit}_{i}^{\ell,h}(\mathbf{a}_{i}^{\ell,h})=W_{U}\sigma(\mathbf{a}_{i}^{\ell,h}),\text{logit}_{i}^{\ell}(\mathbf{m}_{i}^{\ell})=W_{U}\sigma(\mathbf{m}_{i}^{\ell}),\text{logit}_{i}^{\ell}(\mathbf{x}_{i}^{\ell})=W_{U}\sigma(\mathbf{x}_{i}^{\ell})$. Thus, logit lens serves as a powerful tool for pinpointing specific model components crucial to prediction behaviours.
JSD for Context Attribution.
@@
JSD is a symmetrised, smoothed variant of Kullback–Leibler (KL) divergence that quantifies information gap (in bits) between two probability distributions. Because it is symmetric, finite, scale-free, and bounded in $0,\log 2]$, JSD allows scores from different layers to be compared directly, without sensitivity to arbitrary logit shifts. Following “logit-lens” perspective of sun2025redeep , we treat JSD as model’s belief of how much its next-token distribution will change. Concretely, we compute JSD between the full-context token distribution and the distribution obtained after removing a single retrieved sentence $c_{i}$. A high divergence indicates that the model’s internal representations (and therefore its output logits) depend strongly on $c_{i}$. Empirically, ablating the sentence with the highest JSD causes the largest drop in answer likelihood, validating JSD as a concise and reliable signal for context attribution in RAG models (See § 8 for comparisons among JSD, Wasserstein, Total Variation (TV) and Maximum Mean Discrepancy (MMD)).
@@ 4.1 Identifying Relevant Context via JSD
Following assumption proposed by cohen2024contextcite , removal of context segments critical to generating a specific response $\mathcal{R}$ significantly impacts probability distribution of that response. Conversely, the removal of less relevant context segments is expected to minimally affect probability distribution of $\mathcal{R}$.
Unlike the approach by cohen2024contextcite , which requires extensive sampling of ablated contexts for each $(\mathcal{C},\mathcal{Q})$ pair and training a surrogate model to learn context-response relationships, our proposed ARC-JSD method relies purely on inference in the Fig. 1. Specifically, we compute the JSD between the response probability distributions conditioned on the full context $\mathcal{C}$ and on each context-ablated variant $\mathcal{C}_{\text{ABLATE}}(c_{i})$:
@@ 4.2 Evaluation of Context Attribution Accuracy
ARC-JSD (key-value caching) $2PT|\mathcal{C}|^{2}$ 1
To assess efficacy of our ARC-JSD method, we conduct experiments on three widely recognised question-answering datasets commonly used in RAG studies: TyDi QA (clark2020tydi, ): a multilingual QA dataset using entire Wikipedia articles as external context (we only use English part), Hotpot QA (yang2018hotpotqa, ): a multi-hop QA dataset requiring reasoning for questions based on multiple documents, and MuSiQue (trivedi2022musique, ): a high-quality multi-hop QA benchmark over Wikipedia that highlights minimal context and multiple valid reasoning paths to evaluate complex reasoning capabilities. Moreover, we evaluate our ARC-JSD with different training-free baselines for context attribution: ALTI-Logit (ferrando2023explaining, ): a method to directly compare logit difference between input context and generation on token level by accumulating layerwise logit; MIRAGE (qi2024model, ): a gradient-based and token-level method to locate context-sensitive tokens using contrastive feature attribution; Contextcite (cohen2024contextcite, ): a post-hoc method to train a linear surrogate model based on a fixed group of context ablation forward runs. Table 2 summarises the statistics of these datasets, where MuSiQue has the longest context input compared to others with the average length of context in sentences $|\mathcal{C}|=93.6$. Our evaluations involve four instruction-tuned LLMs of varying scales, namely Qwen2-1.5B-IT, Qwen2-7B-IT (yang2024qwen2, ), Gemma2-2B-IT, and Gemma2-9B-IT (team2024gemma, ). For each dataset, we randomly select up to 1,000 samples from their development sets. All models are evaluated in inference mode without further fine-tuning. We mainly evaluate the top-1 context attribution accuracy, which indicates the percentage of overlap between the predicted top-1 context sentence and gold-standard sentence on the datasets [^3] We choose sentence level because current QA datasets only have sentence-level gold labels to evaluate attribution accuracy. However, our ARC-JSD method can be extended to finer-grained interactions such as phrases or sub-sentences spans by dynamically selecting the start and end token indices.. For ALTI-Logit and MIRAGE, which mainly focus on token-level attribution, we use the accumulated operations to locate sentence-level context attribution prediction (Appendix E includes more details).
@@
Table 1 lists theoretical floating-point operations (FLOPs) for each method, where we follow the assumption from kaplan2020scaling ; hoffmann2022training , i.e., one forward pass needs approximately $2PT$ FLOPs. ARC-JSD is considerably cheaper than baselines because it pinpoints salient context sentences without back-propagation or iterative token masking. ContextCite requires a fixed 32 or 256 forward passes; this is economical only when input contains more than 32 or far more than 256 sentences, respectively, but its context attribution accuracy remains below that of ARC-JSD (see Fig. 2(a)). Fig. 2(a) presents compute-accuracy trade-off on MuSiQue dataset across all baselines and LLM backbones. It clearly demonstrates that ARC-JSD consistently outperforms all baselines, yielding an average context attribution accuracy improvement of approximately 10.7%. Although Contextcite-32 is more efficient when $|\mathcal{C}|$ is larger than 32, its attribution accuracy lags behind ARC-JSD. Overall, our method offers substantial computational efficiency improvements, achieving up to 3-fold speedups and consistently align with Pareto-optimal over multiple orders of magnitude for different LLM backbones. In addition, we utilise GPT-4.1 mini as a judge to compare whether generated responses of all RAG models are semantically equivalent to the corresponding gold answers from datasets when context attribution is correct. The average accuracy is up to 99.3% (See Appendix G and F for details).

### v5

# Attributing Response to Context: A Jensen–Shannon Divergence Driven Mechanistic Study of Context Attribution in Retrieval-Augmented Generation

@@ 1 Introduction
Retrieval-Augmented Generation (RAG), leveraging LLMs, has demonstrated significant potential in both academic research (Qian et al., 2024; Yue et al., 2025; Song et al., 2025) and industrial applications (Yang et al., 2024; Wang et al., 2025a; Guo et al., 2025) by enhancing the accuracy and grounding of generated responses through external contexts such as provided documents or retrieved articles online. A key benefit of RAG lies in its ability to mitigate the hallucination by explicitly attributing generated responses to specific segments of the provided context, known as context attribution [^1] We use the term context attribution in this work, and there are several different terms used in this area, such as citation, self-citation, etc. (Wang et al., 2024; Qi et al., 2024; Cohen-Wang et al., 2024; Chuang et al., 2025).
@@
Nevertheless, verifying the extent to which generated responses are genuinely grounded in their cited context remains a challenging task. Current approaches frequently rely heavily on human annotation (Zeng et al., 2021; Menick et al., 2022; Slobodkin et al., 2024) or computationally expensive methods such as model fine-tuning and gradient-based feature attribution for accurate attribution (Yue et al., 2023; Qi et al., 2024; Chuang et al., 2025), particularly when dealing with extensive documents. For instance, Qi et al. (2024) utilised distribution shifts between responses generated with and without context to identify relevant tokens and employed gradient-based feature attribution to pinpoint context relevance. Similarly, Chuang et al. (2025) enhanced context attribution accuracy through reward-driven fine-tuning within a Direct Preference Optimisation (DPO) framework, based on probability drop and hold analysis of model outputs to context ablation.
@@
To circumvent these computationally intensive methods, Cohen-Wang et al. (2024) introduced an inference-time attribution mechanism premised on the assumption that if removing grounded context segments substantially reduces the probability of a generated response, those segments are deemed necessary. Conversely, if retaining only grounded segments maintains response probability, these segments are considered sufficient. By capturing hundreds of probability ablation variations per context-response pair, they trained a linear surrogate model based on those hundreds of vectors, including the context segment masks and the corresponding generation probability of the original response, to identify context segments crucial for grounding model responses.
@@
However, Cohen-Wang et al. (2024) still needs hundreds of RAG model’s forward calls to collect probability ablation samples for the linear surrogate model training. We propose a novel inference-time Jensen–Shannon Divergence driven method to Attribute Response to Context (ARC-JSD), building upon the inference-attribution assumption above. Our method evaluates the divergence in response distributions generated under the full context compared to sentence-ablated contexts, ranking context sentences based on their JSD differences because of JSD’s symmetric, finite, scale-free, and bounded properties (see § 3 and § 8 for details). This approach offers a significant computational advantage, as it eliminates the need for any additional fine-tuning or surrogate modelling. Furthermore, our ARC-JSD can avoid missing or smoothing non-linearities using JSD to directly quantify actual output distribution shift compared to the linear surrogate modelling (Cohen-Wang et al., 2024).
@@
We empirically evaluate our JSD-driven context attribution approach across multiple question-answering benchmarks, i.e., TyDi QA (Clark et al., 2020), Hotpot QA (Yang et al., 2018), and MuSiQue (Trivedi et al., 2022), using state-of-the-art instruction-tuned LLMs including Qwen2-1.5B-Instruct, Qwen2-7B-Instruct (Yang et al., 2024), Gemma2-2B-Instruct, and Gemma2-9B-Instruct (Team et al., 2024). Our results not only demonstrate improved average accuracy over 10% in context attribution but also achieve computational efficiency, achieving up to a three-fold speedup compared to Cohen-Wang et al. (2024)’s linear-surrogate-based and other gradient-based baselines.
@@
Moreover, we investigate deeper into a mechanistic exploration of context attribution within RAGs by integrating JSD-based analysis with Logit Lens (nostalgebraist, 2020). Through systematic probing, we identify specific attention heads and MLP layers critical for context attribution. By subsequently manipulating these located attention heads and MLPs using JSD scores as confidence gates, we can control RAGs and further mitigate the hallucination rate. In addition, we can visualise how relevant knowledge is stored in the corresponding MLP layers.
@@ 2 Related Work
Context attribution for RAG. Prior works for context attribution mainly focus on teaching RAG LLMs to generate self-citations for responses, such as few-shot in-context learning (Gao et al., 2023), instruction fine-tuning (Ye et al., 2024). Some post-hoc works (Chen et al., 2023; Qi et al., 2024) used an auxiliary language model or gradient-based feature attribution to locate relevant context segments. In general, those methods for context attribution are corroborative (Worledge et al., 2024) in nature, as citations within context are evaluated on whether they support or imply a generated response. Meanwhile, Cohen-Wang et al. (2024); Chuang et al. (2025); Cohen-Wang et al. (2025); Liu et al. (2024) including our work focus on the contributive attribution methods, which are used to identify whether citations cause RAG LLMs to generate a response. Chuang et al. (2025) proposed a reward-based fine-tuning with DPO to guide RAG LLMs for context attribution, and Cohen-Wang et al. (2024; 2025) further trained a linear surrogate model to identify context segments crucial for grounding model responses. Liu et al. (2024) focuses on formalising and comparing different attribution acceleration methods and ignores attribution accuracy improvements. However, compared to Cohen-Wang et al. (2024); Chuang et al. (2025); Cohen-Wang et al. (2025) and corroborative methods above, our ARC-JSD method eliminates the need for any additional fine-tuning or surrogate modelling, and it can be directly integrated into any existing RAGs.
@@
Mechanistic analysis for RAG. Existing mechanistic studies focus on next token generation task to analyse internal mechanisms of attention heads or MLPs, such as hallucination detection (Ferrando et al., 2025), multiple-choice questions (Li and Gao, 2024; Wiegreffe et al., 2025; Wang et al., 2025b) and knowledge editing (Meng et al., 2022a; 2023; Katz et al., 2024). Recently, Sun et al. (2025) used a mechanistic interpretability method to analyse attention heads and MLPs of RAGs for the hallucination detection task. Compared to Sun et al. (2025) focusing on locating sources which lead to hallucinations, our proposed ARC-JSD can be regarded as a complementary method to locate citations within context segments and analyse attentions and MLPs, which causes RAG LLMs to generate a correct response. Wu et al. (2025a) focuses on mechanistically analysing retrieval attention heads of RAG LLMs under the Needle-in-the-Haystack (NIAH) setting, where they mainly evaluate whether retrieval attention heads conduct a copy-and-paste operation for retrieving a semantically irrelevant “needle” sentence from the context to the model’s outputs. Compared to Wu et al. (2025a), which restricts their mechanistic analysis to the NIAH setting where the model performs copy-and-paste retrieval, our work investigates how RAG LLMs mechanistically generate responses based on retrieved content through paraphrasing and contextual integration. This setting better reflects real-world RAG applications [^2] Compared to traditional RAG to directly map prompts based on their word embeddings, our work has a more general setting, which avoids potential embedding mismatch due to the common paraphrase of RAGs., where models rarely copy text exactly but instead synthesise and rephrase information from retrieved sources.
@@ 3 Background
Problem Setup. Consider an autoregressive Transformer-based language model (LLM), denoted as $\mathcal{P}_{\text{LM}}(\cdot)$. Under RAG settings, this model generates responses ($\mathcal{R}$) based on an input query ($\mathcal{Q}$) and associated context ($\mathcal{C}$). Formally, response generation process can be described as $\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|\mathcal{C},\mathcal{Q})$, where context $\mathcal{C}$ consists of sentences $(c_{1},c_{2},\dots,c_{|\mathcal{C}|})$, the query $\mathcal{Q}$ comprises tokens $(q_{1},q_{2},\dots,q_{|\mathcal{Q}|})$, and the generated response $\mathcal{R}$ includes tokens $(r_{1},r_{2},\dots,r_{|\mathcal{R}|})$. Our analysis of context attribution focuses on how the entire response distribution changes when conditioned on the full context set and ablated context alongside the query: $\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|c_{1},\dots,c_{|\mathcal{C}|},\mathcal{Q}),\>\mathcal{R}\sim\mathcal{P}_{\text{LM}}(\cdot|\mathcal{C}_{\text{ABLATE}}(c_{i}),\mathcal{Q})\,\text{where}\,\mathcal{C}_{\text{ABLATE}}(c_{i})=\mathcal{C}\setminus\{c_{i}\},\>i\in\{1,\dots,|\mathcal{C}|\}$.
@@
Logit Lens. Logit lens (nostalgebraist, 2020) is a mechanistic interpretability method designed to analyse intermediate representations within autoregressive Transformers. Given the LLM architecture described in Appendix D, logit lens leverages intermediate representations to quantify the direct contribution of attention heads ($\mathbf{a}_{i}^{\ell,h}$), MLP outputs ($\mathbf{m}_{i}^{\ell}$), and residual streams ($\mathbf{x}_{i}^{\ell}$) to token logits: $\text{logit}_{i}^{\ell,h}(\mathbf{a}_{i}^{\ell,h})=W_{U}\sigma(\mathbf{a}_{i}^{\ell,h}),\text{logit}_{i}^{\ell}(\mathbf{m}_{i}^{\ell})=W_{U}\sigma(\mathbf{m}_{i}^{\ell}),\text{logit}_{i}^{\ell}(\mathbf{x}_{i}^{\ell})=W_{U}\sigma(\mathbf{x}_{i}^{\ell})$. Thus, logit lens serves as a powerful tool for pinpointing specific model components crucial to prediction behaviours.
@@
JSD for Context Attribution. JSD is a symmetrised, smoothed variant of Kullback–Leibler (KL) divergence that quantifies information gap (in bits) between two probability distributions. Because it is symmetric, finite, scale-free, and bounded in $0,\log 2]$, JSD allows scores from different layers to be compared directly, without sensitivity to arbitrary logit shifts. Following “logit lens” perspective of Sun et al. (2025), we treat JSD as model’s belief of how much its next-token distribution will change. Concretely, we compute JSD between the full-context token distribution and the distribution obtained after removing a single retrieved sentence $c_{i}$. A high divergence indicates that the model’s internal representations (and therefore its output logits) depend strongly on $c_{i}$. Empirically, ablating the sentence with the highest JSD causes the largest drop in answer likelihood, validating JSD as a concise and reliable signal for context attribution in RAG models (See § 8 for comparisons among JSD, Wasserstein, Total Variation (TV) and Maximum Mean Discrepancy (MMD)).
@@ 4.1 Identifying Relevant Context via JSD
Following assumption proposed by Cohen-Wang et al. (2024), removal of context segments critical to generating a specific response $\mathcal{R}$ significantly impacts the probability distribution of that response. Conversely, the removal of less relevant context segments is expected to minimally affect $\mathcal{R}$.
Unlike the approach by Cohen-Wang et al. (2024), which requires extensive sampling of ablated contexts for each $(\mathcal{C},\mathcal{Q})$ pair and training a surrogate model to learn context-response relationships, our proposed ARC-JSD method relies purely on inference in the Fig. 1. Specifically, we compute the JSD between the response probability distributions conditioned on the full context $\mathcal{C}$ and on each context-ablated variant $\mathcal{C}_{\text{ABLATE}}(c_{i})$:
@@ 4.2 Evaluation of Context Attribution Accuracy
ARC-JSD $2PT|\mathcal{C}|^{2}$ 1
To assess efficacy of our ARC-JSD method, we conduct experiments on three widely recognised question-answering datasets commonly used in RAG studies: TyDi QA (Clark et al., 2020): a multilingual QA dataset using entire Wikipedia articles as external context (we only use English part), Hotpot QA (Yang et al., 2018): a multi-hop QA dataset requiring reasoning for questions based on multiple documents, and MuSiQue (Trivedi et al., 2022): a high-quality multi-hop QA benchmark over Wikipedia that highlights minimal context and multiple valid reasoning paths to evaluate complex reasoning capabilities. Moreover, we evaluate our ARC-JSD with different training-free baselines for context attribution: ALTI-Logit (Ferrando et al., 2023): a method to directly compare logit difference between input context and generation on token level by accumulating layerwise logit; MIRAGE (Qi et al., 2024): a gradient-based and token-level method to locate context-sensitive tokens using contrastive feature attribution; Contextcite (Cohen-Wang et al., 2024): a post-hoc method to train a linear surrogate model based on a fixed group of context ablation forward runs. Table 2 summarises the statistics of these datasets, where MuSiQue has the longest context input compared to others, with the average length of context in sentences $|\mathcal{C}|=93.6$. Our evaluations involve four instruction-tuned LLMs of varying scales, namely Qwen2-1.5B-IT, Qwen2-7B-IT (Yang et al., 2024), Gemma2-2B-IT, and Gemma2-9B-IT (Team et al., 2024). For each dataset, we randomly select up to 1,000 samples from its development sets. All models are evaluated in inference mode without further fine-tuning. We mainly evaluate the top-1 context attribution accuracy, which indicates the percentage of overlap between the predicted top-1 context sentence and gold-standard sentence on the datasets [^3] We choose sentence level because current QA datasets only have sentence-level gold labels to evaluate attribution accuracy. However, our ARC-JSD method can be extended to finer-grained interactions such as phrases or sub-sentence spans by dynamically selecting the start and end token indices.. For ALTI-Logit and MIRAGE, which mainly focus on token-level attribution, we use the accumulated operations to locate sentence-level context attribution prediction (Appendix E includes more details).
@@
Table 1 lists theoretical floating-point operations (FLOPs) for each method, where we follow the assumption from Kaplan et al. (2020); Hoffmann et al. (2022), i.e., one forward pass needs approximately $2PT$ FLOPs. ARC-JSD is considerably cheaper than baselines because it pinpoints salient context sentences without back-propagation or iterative token masking. ContextCite requires a fixed 32 or 256 forward passes; this is economical only when input contains more than 32 or far more than 256 sentences, respectively, but its context attribution accuracy remains below that of ARC-JSD (see Fig. 2(a)). Fig. 2(a) presents compute-accuracy trade-off on MuSiQue dataset across all baselines and LLM backbones. It clearly demonstrates that ARC-JSD consistently outperforms all baselines, yielding an average context attribution accuracy improvement of approximately 10.7%. Although Contextcite-32 is more efficient when $|\mathcal{C}|$ is larger than 32, its attribution accuracy lags behind ARC-JSD. Overall, our method offers substantial computational efficiency improvements, achieving up to 3-fold speedups and consistently aligns with Pareto-optimal over multiple orders of magnitude for different LLM backbones. In addition, we utilise GPT-4.1 mini as a judge to compare whether generated responses of all RAG models are semantically equivalent to the corresponding gold answers from datasets when context attribution is correct. The average accuracy is up to 99.3% (See Appendix H and F for details).